1.      A cannonball is catapulated toward a castle. The cannon ball velocity when it leaves a catapult is 40 m/s at an angle of 370 with respect to the horizontal and the cannonball is 7.0 m from the ground at this time. (a) What is the maximum height above the ground reach by the cannon ball? (b) What horizontal Distance from the release point will it land?


1
Expert's answer
2020-08-31T12:47:42-0400

The maximum height is


"H=h+\frac{(v\text{ sin}\alpha)^2}{2g}=7+\frac{(40\text{ sin}37°)^2}{2\cdot9.8}=37\text{ m}."

The horizontal distance is determined by the horizontal component of initial velocity and time of flight:


"R=v\text{ cos}37°t."

The time of flight is made up of time t1 required to go up from 7-m level, and then t2 -from max. height to the ground:

"t_1=\frac{v\text{ sin}\alpha}{g}=2.46\text{ s},\\\space\\
t_2=\sqrt{\frac{2H}{g}}=2.75\text{ s},\\\space\\
R=166\text{ m}."